OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9(F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (anonymous):

I have to answer hold on. i did the same thing

OpenStudy (anonymous):

okay do you need the work also?

OpenStudy (anonymous):

wow i passed high school and i dont know this

OpenStudy (anonymous):

aditi

OpenStudy (anonymous):

Part 1: Solve the formula for F. Show all steps. C = 5/9 (F – 32) 9C = 5(F - 32) F - 32 = 9C/5 F = 9C/5 + 32 -------------- Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is. F = 9*6/5 + 32 F = 54/5 + 32 F = 10.8 + 32 F = 42.8

OpenStudy (anonymous):

DONE :)

OpenStudy (anonymous):

yea i looked it up aswell so thats what i found to.

OpenStudy (anonymous):

@macmiller101 i found this on a website too. but thanks for copying it for me to make sure

OpenStudy (anonymous):

welcome lol